solve by the elimination method p=40+q and 7p=-9q+360
Found 2 solutions by stanbon, Menjax:
Answer by stanbon(75887)

solve by the elimination method
p=40+q
7p=-9q+360
--------
Rearrange:
p - q = 40
7p + 9q = 360
-------------------
Modify for elimination::
7p - 7q = 7*40
7p + 9q = 360
---------------------
Subtract and solve for "q"::
16q = 80
----
q = 5
-----
Solve for "p"::
p - q = 40
p - 5 = 40
p = 45
=====================
Cheers,
Stan H.
==============

Answer by Menjax(62)

p=40+q and 7p=-9q+360
so we will times p=40+q by 7,
7p=280+7q
now we have
7p=280+7q
7p=360-9q
we will minus the first equation from the second
we are left with
0=80-16q
16q=80
q=5
put q is 5 into 7p=280+7q to find p
7p = 280 + 35
p=315/7
p=45 and q=5